OpenStudy (anonymous):

Find the quadratic equation with roots -1+ 4i and -1 - 4i

OpenStudy (anonymous):

just write it in the form (x-a)(x-b)=0 and expand the brackets, where a and b are the given roots

OpenStudy (anonymous):

so (4-1) (-4-1) ?

OpenStudy (anonymous):

not sure what u mean - the roots are complex numbers

OpenStudy (anonymous):

do you know the answer? maybe ill kind of get it more idk

OpenStudy (anonymous):

have you met complex numbers before ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

you really need to know some complex algebra to tackle this question the i that appears in your question is the imaginary unit, defined so that i squared = -1 that's about all you need to know get the answer here

OpenStudy (anonymous):

4i = -1 ?

OpenStudy (anonymous):

im very bad at this erh

OpenStudy (anonymous):

you need to read up on quadratic equations a bit

OpenStudy (anonymous):

i really should

OpenStudy (anonymous):

but could you show me how to do this step by step or nah...

OpenStudy (anonymous):

do you understand what the root of a quadratic equation is ?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

ok, and any quadratic has 2 roots

OpenStudy (anonymous):

if you call the roots a and b, then you can reconstruct the original equation by writing it as (x-a)(x-b)=0 and expanding the brackets

OpenStudy (anonymous):

now in your particular example, a = -1+4i and b = -1-4i

OpenStudy (anonymous):

so you have to expand (x+1-4i)(x+1+4i) = 0

OpenStudy (anonymous):

just multiply out the brackets and remember that wherever you get i^2, you can replace it with -1